OpenStudy (4everaddicted2anime):

An Native American tepee is a conical tent. Find the number of cubic feet of air in a teepee 10 ft. in diameter and 12 ft. high. Answer must be in terms of pi

OpenStudy (welshfella):

volume of a cone = (1/3) pi r^2 h where r = radius of the base , h = height ( note r = 1/2 * diameter).

OpenStudy (welshfella):

No its r^2 not r

OpenStudy (4everaddicted2anime):

So it would be (1/3)pi (5)^2 * 12 Then I get 1/3 pi 25 * 12 = (1/3) * 300pi So the answer would be 100pi right?

OpenStudy (welshfella):

yes

OpenStudy (4everaddicted2anime):

Okay. Thank you

OpenStudy (welshfella):

100 pi cu ft

OpenStudy (welshfella):

yw